-2(x+5)=30
4x+10=30
4x+10=30+40
4x +10=70
I know this is wrong but where did I go wrong?

Answer by ilana(305)

I see a few errors in your process. First of all, I am not sure what made you say -2(x+5)=4x+10. When you distribute the -2, you should get -2x-10. So now we have -2x-10=30. Now, we want the terms with x alone on 1 side, so we need to move the 10. The 10 is being subtracted, so we add 10 to both sides. So -2x-10+10=30+10, so -2x=40. Now we need just x, so divide by -2 and x=-20.